OpenStudy (anonymous):

is x+x+x+2x 2x? or is it 2x^3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

i just need to know if x+x+x+2x = 2x or 2x^3

OpenStudy (skullpatrol):

What do you think it is?

OpenStudy (skullpatrol):

And why?

OpenStudy (yrelhan4):

look. x+x is 2x x*x is x^2.. that should be enough info for your question.

OpenStudy (anonymous):

okay so its 2x^3 ? thanks!

OpenStudy (yrelhan4):

eh. no it isnt. i dont see the multiplication sign? where did the x^3 part come from? you only have the plus signs right? unless you mistyped the question.

OpenStudy (anonymous):

yeah its only plus signs

OpenStudy (anonymous):

so its just 2x

OpenStudy (skullpatrol):

@allie22 you are not explaining why you think that way?

OpenStudy (yrelhan4):

no.. there are 3 x's and one 2x to be added.

OpenStudy (anonymous):

because at first i thought it was just 2x sense it was addition and x+x is x, but then i second guessed myself and was like wait would it be 2x^3 and now idk so yeah

OpenStudy (skullpatrol):

Recall the distributive property ab + ac = a(b +c)

OpenStudy (skullpatrol):

Now look at this equation x+x+x+2x = x(1 + 1 + 1 + 2)

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

then you would have 1x + 1x + 1x + 2 correct?

OpenStudy (anonymous):

+2x*

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

so now what

OpenStudy (anonymous):

3x +2x right

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

so thats the final answer?

OpenStudy (skullpatrol):

Keep going you can go further.

OpenStudy (anonymous):

wait how because you cant add like terms

OpenStudy (anonymous):

you can only add like terms*

OpenStudy (skullpatrol):

3x +2x = x(3 + 2)

OpenStudy (anonymous):

thats still 3x+2x .. so now what

OpenStudy (skullpatrol):

$$\Huge ab + ac = a(b +c)$$

OpenStudy (anonymous):

i dont get what your trying to say how do i simplify 3x+2x

OpenStudy (skullpatrol):

$$\Huge 3x +2x = x(3 + 2)$$

OpenStudy (anonymous):

omg. i dont get it.

OpenStudy (skullpatrol):

What is 3 + 2 = ?

OpenStudy (anonymous):

5.

OpenStudy (anonymous):

but you cant add 3x + 2x i thought because there not like terms

OpenStudy (skullpatrol):

What is the definition of "like terms"?